# M2SE: A Multistage Multitask Instruction Tuning Strategy for Unified Sentiment and Emotion Analysis

### v2

@@ IV-A Experiments Setup
IV-A3 Emotion Cause-pair Extraction.
@@
IV-B Implementation Details
For the visual encoder, we utilize pre-train ViT with input images resized to $448\times 448$ pixels. During the fine-tuning process, the visual backbone is frozen, with the focus placed on training the linear projection layer. For LLM, we employ Phi-3-Mini[45] for EmoVerse-4B and InternLM-2.5-7B-Chat [^3] https://github.com/InternLM/InternLMfor EmoVerse-8B, both augmented with LoRA[46] for parameter-efficient fine-tuning.
In each stage of training, we train for 2 epochs with LoRA hyperparameters set to $r=8$ and $a=32$. The learning rate is set to $1\times 10^{-5}$, and we use a warm-up cosine schedule for the learning rate decay. For EmoVerse-4B, the training is conducted on 2 NVIDIA RTX 3090 GPUs, requiring approximately 48 hours. For EmoVerse-8B, the training is executed on a NVIDIA A100 GPU, taking around 55 hours.
IV-C Results
IV-C1 Performance comparison.
For MSA task, we compare our EmoVerse model with several SOTA baselines, including TFN[7], MulT[12], PMR[21], MISA[17], MMIM[14], Self-MM [13], BBFM[15], FDMER[16],CONKI[18], DMD[20], ConFEDE[19], and MRC-D3AE[27]. In ER task, we evaluate our model against baseline models with DialogueGCN [22], DialogueCRN[23], DAG-ERC[8], MM-DFN[24], EmoCaps[9], ERMC-DisGCN[25], MMGCN[26], GA2MIF[10], UniMSE[11], and FacialMMT[47].
Additionally, for both MSA and ER, we select advanced models capable of processing both images and videos. These include LLaVA-Video [48] and InternVL2 [49], both of which are designed to handle multimodal input efficiently. The performance comparison results, presented in Table III and Table IV, clearly demonstrate that EmoVerse achieves SOTA performance across both tasks, surpassing all the aforementioned baseline models in terms of accuracy and robustness.
For the ECPE task, we conduct tests solely on MLLMs, as they are better equipped to handle the complexity and multimodal nature of the task. The Table V shows the results of ECPE.
TABLE III: The performance comparison on CMU-MOSEI dataset.
Model Acc2 TFN - / 82.5 MulT - / 82.5 Self-MM 82.81 / 85.17 MMIM 82.24 / 85.97 BBFN - / 86.20 FDMER - / 86.10 MISA 83.60 / 85.50 ConKI 82.73 / 86.25 ConFEDE 81.65 / 85.82 DMD - / 86.60 PMR - / 83.30 MRC-D3AE 83.10 / 85.50 LLaVA-Video - / 82.82 InternVL2-4B - / 82.09 InternVL2-8B - / 85.30 EmoVerse-4B 85.51 / 87.80 EmoVerse-8B 85.93 / 88.51
TABLE IV: The performance comparison on MELD dataset.
Model Acc W.F1-Score DialogueGCN 59.46 58.10 DialogueCRN 60.73 58.39 DAG-ERC - 63.65 MM-DFN - 59.46 EmoCaps - 64.00 ERMC-DisGCN - 64.22 MMGCN 62.49 58.65 GA2MIF - 58.94 UniMSE 65.09 65.51 FacialMMT - 66.58 LLaVA-Video 23.87 22.65 InternVL2-4B 39.46 40.59 InternVL2-8B 42.18 44.13 EmoVerse-4B 67.28 66.58 EmoVerse-8B 67.78 66.74
TABLE V: The performance comparison on ECF2 dataset.
Model W.F1-Score F1-Score LLaVA-Video 22.59 22.78 InternVL2-4B 33.24 33.35 InternVL2-8B 40.57 40.51 EmoVerse-4B 72.30 72.21 EmoVerse-8B 73.54 73.62
IV-C2 Analysis of Emotion Reasoning.
To demonstrate the qualitative performance of EmoVerse, we present a detailed comparison of emotion reasoning results across various models. Figure 4 showcases the emotion reasoning outputs of five models: GPT-4o, Video-LLaMA2[50], LLaMA-Video, InternVL2, and EmoVerse. The video depicts a woman displaying a strong reaction of surprise, highlighting the models’ ability to infer emotional states from multimodal inputs.
GPT-4o and Video-LLaMA2 both classify the emotion in the video as ”confident” or ”assertive.” However, LLaMA-Video and InternVL2 correctly identify the emotion as ”surprise.” Despite this, these two models fail to fully integrate multimodal information for ERI. For instance, LLaMA-Video makes its inference solely based on facial expressions, while InternVL2 briefly acknowledges the facial expressions without providing a detailed explanation of the emotional cause, instead relying solely on textual cues for emotion reasoning. In contrast, EmoVerse effectively combines both multimodal video and text information to infer the cause of the emotion, yielding a correct and comprehensive emotional interpretation.
@@ IV-C Results
Figure 4: Comparison of different models on ERI. It can be seen that both GPT-4o and Video-LLaMA2 make errors in ER. While Llava-Video and InternVL2 correctly classify the emotions, their descriptions of the reasons behind the emotions are incomplete. In contrast, EmoVerse not only correctly classifies the emotions but also provides a comprehensive description of the reasons behind them.
IV-D Ablation Evaluation
We use EmoVerse-4B to do the following ablation experiments.
IV-D1 Comparison of the M2SE Strategy with Single-Task Fine-Tuning.
To evaluate the impact of the M2SE strategy within the same model framework and compare its performance across various tasks, we conduct an ablation study. As shown in Table VI, the tasks in parentheses represent the fine-tuning tasks, with each task corresponding to the same datasets used in our evaluation. ”Z.S.” stands for zero-shot. The M2SE strategy is trained using the EMT dataset. The results demonstrate that, although the performance on some metrics may not exceed that of single-task learning, EmoVerse trained with the M2SE strategy consistently performs better overall compared to other models.
@@ IV-D Ablation Evaluation
TABLE VI: Comparison of the M2SE Strategy with Single-Task Fine-Tuning
Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.28 66.58 72.30 72.21
IV-D2 Effect of the FER Task.
To explore the contribution of the FER task to MSA, we conduct an ablation experiment to compare the effects of increasing and decreasing the FER sampling ratio on MSA performance, using the same training data volume in the first stage. For a low sampling ratio, we denote it as ”lf,” and for a high sampling ratio, we denote it as ”hf.” As shown in Table VII, under the same data volume, increasing the FER task ratio—despite the model performing fewer MSA tasks—still leads to stronger learning performance. This indicates that even with fewer MSA samples, the contribution of FER significantly enhances the model’s ability to perform sentiment analysis.
TABLE VII: Effect of the FER Task.
Model Acc2 EmoVerse (lf) 77.09 / 87.47 EmoVerse (hf) 82.61 / 87.50
IV-D3 Effect of the Choice of Tasks in the Second Stage.
To investigate how task configurations impact model performance and better understand cross-task interactions, we conduct an ablation study in the second stage with three setups: ”t1” (only ER task), ”t3” (ER, ERI, and ECPE tasks), and ”t4” (all four tasks). Results, as shown in Table VIII, show that focusing solely on ER (t1) improves ER performance but weakens MSA and emotion reasoning tasks. Adding ERI and ECPE (t3) slightly lowers MSA and ER performance but enhances the model’s emotion reasoning ability. When MSA is reintroduced with a small sampling rate (t4), the model effectively integrates cross-task knowledge, improving performance across all tasks and surpassing first-stage MSA results. This demonstrates the importance of balanced task integration for holistic model improvement.
TABLE VIII: Effect of the Distribution of Tasks on EmoVerse.
Model CMU-MOSEI MELD ECF2.0
Acc2 Acc W.F1 F1 W.F1
EmoVerse (f) 82.61 / 87.50 - - - -
EmoVerse (t1) 81.68 / 87.20 66.74 66.24 25.40 26.69
EmoVerse (t3) 79.0 / 86.78 66.97 65.67 69.85 69.95
EmoVerse (t4) 85.51 / 87.80 67.78 66.74 72.30 72.21
V Conclusion
In this paper, we propose the M2SE strategy to guide the training process, enabling the model to learn from multiple sentiment- and emotion-related tasks across different stages. To support this, we construct the EMT dataset, which includes a large number of annotated samples spanning a wide range of tasks. Additionally, we introduce EmoVerse, a unified MLLM designed to enhance emotion perception by effectively integrating both visual and textual data. By leveraging multitask and multistage learning, EmoVerse unifies sentiment and emotion tasks, achieving superior performance in MSA, ER, and EPCE. It outperforms existing models on benchmark datasets such as CMU-MOSEI and MELD. The experimental results highlight the effectiveness of the M2SE strategy in enabling the model to capture and reason about complex emotional contexts, showcasing its potential to drive advancements in emotion-aware AI applications.

### v3

@@ 5.1.2 Emotion Recognition
5.1.3 Emotion Cause-pair Extraction
@@
Figure 4: Comparison of different models on ERI. It can be seen that both GPT-4o and Video-LLaMA2 make errors in MER. While Llava-Video and InternVL2 correctly classify the emotions, their descriptions of the reasons behind the emotions are incomplete. In contrast, EmoVerse not only correctly classifies the emotions but also provides a comprehensive description of the reasons behind them.
5.2 Implementation Details
For the visual encoder, we use a pre-trained ViT with input images resized to $448\times 448$ pixels. During fine-tuning, the visual backbone is frozen, and training focuses on the linear projection layer. For the LLM, EmoVerse-4B uses Phi-3-Mini Abdin et al. (2024), and EmoVerse-8B uses InternLM-2.5-7B-Chat [^3] https://github.com/InternLM/InternLM, both augmented with LoRA Hu et al. (2021b) for parameter-efficient fine-tuning.
Training is done for 2 epochs with LoRA hyperparameters $r=8$ and $a=32$, a learning rate of $1\times 10^{-5}$, and a warm-up cosine decay schedule. EmoVerse-4B training takes approximately 48 hours on 2 NVIDIA RTX 3090 GPUs, while EmoVerse-8B requires about 55 hours on a single NVIDIA A100 GPU.
5.3 Results
5.3.1 Performance comparison
For MSA task, we compare our EmoVerse model with several SOTA baselines, including MulT Tsai et al. (2019), PMR Lv et al. (2021), MISA Hazarika et al. (2020), MMIM Han et al. (2021b), Self-MM Yu et al. (2021), BBFM Han et al. (2021a), CONKI Yu et al. (2023), ConFEDE Yang et al. (2023), and MRC-D3AE Zhang et al. (2024a). In MER task, we evaluate our model against baseline models with DialogueGCN Ghosal et al. (2019), DialogueCRN Hu et al. (2021a), DAG-ERC Shen et al. (2021), MM-DFN Hu et al. (2022a), EmoCaps Li et al. (2022), DisGCN Sun et al. (2021), MMGCN Hu et al. (2021c), GA2MIF Li et al. (2023a), UniMSE Hu et al. (2022b), and FacialMMT Zheng et al. (2023).
Additionally, for both MSA and MER, we select models capable of processing both images and videos, including LLaVA-VideoZhang et al. (2024b) and InternVL2Chen et al. (2024b). For the ECPE task, we conduct tests solely on MLLMs, as they are better equipped to handle the complexity and multimodal nature of the task. The performance comparison results, shown in Table 4 clearly demonstrate that EmoVerse achieves SOTA performance across both tasks, outperforming all the aforementioned baselines.
5.3.2 Analysis of Emotion Reasoning
To demonstrate EmoVerse’s qualitative performance, we compare emotion reasoning results across five models: GPT-4o, Video-LLaMA2Cheng et al. (2024b), LLaMA-Video, InternVL2, and EmoVerse, as shown in Figure 4. The video depicts a woman displaying a strong surprise reaction, highlighting the models’ ability to infer emotional states from multimodal inputs.
GPT-4o and Video-LLaMA2 misclassify the emotion as ”confident” or ”assertive,” while LLaMA-Video and InternVL2 correctly identify it as ”surprise.” However, these models fail to fully integrate multimodal information for emotion reasoning. LLaMA-Video relies solely on facial expressions, while InternVL2 combines facial cues with textual analysis but lacks detailed explanation of the emotional cause. In contrast, EmoVerse effectively combines video and text to accurately infer the emotional cause, providing a comprehensive interpretation.
Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Zero-shot) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(MER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.78 66.74 72.30 72.21
@@ 5.3.2 Analysis of Emotion Reasoning
Table 5: Comparison of the M2SE Strategy with Single-Task Fine-Tuning
5.4 Comparison of the M2SE Strategy with Single-Task Fine-Tuning
To assess the impact of the M2SE strategy within the same framework and compare its performance across tasks, we conduct an ablation study. Table 5 lists the fine-tuning tasks (in parentheses) with their corresponding datasets. The M2SE strategy uses the AMT dataset for training. Results show that while some metrics may fall short of single-task learning, EmoVerse with M2SE outperforms other models overall.
6 Conclusion
In this work, we explore the impact of different affective tasks on MLLM under a multistage and multitask training framework, identifying and validating the optimal strategy called M2SE. Using the M2SE strategy, we train EmoVerse, addressing the lack of a general-purpose MLLM in affective computing. Additionally, we construct the AMT dataset, designed to support sentiment and emotion perception training for MLLMs. Experimental results demonstrate that EmoVerse achieves state-of-the-art performance across multiple tasks. This study highlights the effectiveness of the M2SE strategy in unifying sentiment and emotion tasks, paving the way for more emotionally intelligent AI systems.